OpenStudy (anonymous):

the endpoints of CD are C(6,1) and D (-4, -1). Find the midpoint M of CD . can someone help me?

OpenStudy (danjs):

The midpoint will be half of the chenge in y value, and half of the change in x value

OpenStudy (danjs):

\[\frac{ y2 + y1 }{ 2 } = \frac{ -1 + 1 }{ 2 } = \frac{ 0 }{ 2 } = 0 \] Do that for the x value now

OpenStudy (danjs):

(x,0) is what you have so far

OpenStudy (anonymous):

so I got (1,0) thanks!

OpenStudy (danjs):

yep, no prob